# Inducing Causal World Models in LLMs for Zero-Shot Physical Reasoning

### v3

@@ V-A2 Performance on the PhysiCa-Bench Benchmark
The FSPA directly measures the fidelity of the CPM as a physics simulator by calculating the Mean Squared Error between its predicted final state and the ground-truth state from the simulation. Our model achieves a low FSPA of 0.08, indicating that its internal world model generates accurate, quantitative predictions. Baseline LLMs cannot be evaluated on this metric as they lack the architectural components to produce a structured state prediction, a key differentiator of our approach. This capacity for state prediction is a foundational element of classic world models in robotics and reinforcement learning [39, 40], and our work successfully integrates this principle into the LLM paradigm.
The most compelling evidence for our framework’s success comes from the Causal Consistency Score (CCS). This strict metric requires the model to correctly answer both a factual question and its corresponding counterfactual counterpart. A high CCS is difficult to achieve through superficial pattern matching; it requires the model to understand the specific causal link between the variable that was changed (the intervention) and the change in the outcome. Our CWMI model achieves a remarkable CCS of 87.6%. In stark contrast, Llama 3 and GPT-4 score only 12.3% and 21.9%, respectively. Their scores, while better than random chance, suggest they are answering the questions independently, with no consistent underlying causal model to link the factual and counterfactual pairs. They may correctly guess that a steel ball falls faster than a rubber one, but they fail to do so with the consistency that would imply genuine understanding. This finding strongly supports our core thesis: the interventional training paradigm described in Section IV (Eq. 6) is exceptionally effective at forcing a model to move beyond correlation and learn the causal structure of the physical world. This aligns with broader calls in the AI community for a more central role for causality in building robust and generalizable intelligence [42, 43].
@@ V-B Ablation Studies: Deconstructing the CWMI Framework
To understand the source of CWMI’s performance and validate our specific design choices, we conducted a series of ablation studies on PhysiCa-Bench. We systematically removed or altered key components of our framework and measured the impact on performance. The results, shown in Table LABEL:table:ablation_study, unequivocally confirm the importance of each part of our methodology.
@@ V-B1 The Critical Role of the Causal Intervention Loss
As Table LABEL:table:ablation_study shows, removing $L_{causal}$ leads to a catastrophic drop in causal reasoning ability. While the model still achieves a respectable accuracy of 85.3% and a low FSPA of 0.11 (proving it can learn to predict outcomes from observation), its Causal Consistency Score plummets from 87.6% to 43.2%. This is the most important result of our ablation analysis. It provides direct evidence that a model trained purely on observational data, even with a perfect world model architecture, is not guaranteed to learn the correct causal dependencies. It may learn spurious correlations present in the data—for instance, if heavier objects in the training set also tend to be darker, the model might incorrectly associate color with gravitational acceleration. The $L_{causal}$ term breaks these correlations by forcing the model to isolate the effect of a single, specific intervention. This result empirically validates theoretical arguments about the necessity of interventional data for causal discovery and demonstrates how these principles can be practically applied to train large-scale neural models.
@@ V-B3 Validating the Architectural Separation
Finally, to justify our two-component architecture, we conducted an experiment where we discarded the CPM entirely and instead fine-tuned the end layers of the Llama 3 LLM directly on the physical reasoning tasks. This ”LLM Fine-tuning (No CPM)” approach is representative of standard domain adaptation techniques. Its performance is poor across the board, barely improving upon the zero-shot GPT-4 baseline and showing almost no causal consistency. This result strongly validates our architectural decision to separate linguistic parsing from physical simulation. A standard Transformer architecture, while a universal approximator in theory, lacks the specific inductive biases for modeling physical dynamics efficiently. Our CPM, while also Transformer-based, is specialized for this role. It operates on a structured state representation and is trained with a physics-oriented objective, making it far more effective and parameter-efficient for this class of problems. This finding echoes the principles of modularity and relational inductive biases [41, 57], suggesting that building complex AI systems may require composing specialized modules rather than relying on a single, monolithic model for all tasks.
@@ V-C1 The Effect of Causal Physics Module Capacity
To investigate the relationship between the capacity of our world model and its performance, we trained several versions of CWMI where we varied the size of the Causal Physics Module (from 64M to 512M parameters). The results, plotted in Fig. LABEL:fig:linechart_scaling, show a clear and positive trend. As the number of parameters in the CPM increases, the performance on PhysiCa-Bench, particularly the FSPA and CCS metrics, steadily improves before beginning to plateau around the 256M parameter mark.
@@
This trend is highly informative. The initial sharp increase in performance suggests that a certain model capacity is required to capture the fundamental physical laws present in our dataset. As the model grows, it becomes more adept at modeling complex interactions and second-order effects. The eventual plateau indicates that, for the complexity of physics represented in PhysiCa-Bench (primarily rigid-body dynamics), our 256M-parameter CPM is near-optimal. This scaling analysis provides confidence in our model design and suggests a clear path for future work: to tackle more complex physical domains like fluid or non-rigid dynamics, a correspondingly larger and perhaps architecturally different CPM will be required. This mirrors the scaling laws that have been so fundamental to the progress of LLMs themselves [2].
@@ V-C2 Qualitative Analysis: When It Works and When It Fails
These limitations are not surprising; they define the boundaries of our current contribution and illuminate the path forward. Addressing them will likely require architectural enhancements to the CPM, such as adopting graph neural networks for multi-object tracking [41], incorporating physics-informed neural network techniques for continuous fields [49], or integrating more sophisticated 3D visual representations [61]. Furthermore, creating more robust systems that can handle noisy or incomplete descriptions is a shared challenge across many AI domains, from facial expression recognition from heterogeneous sources [60] to building reliable WiFi-based sensing systems in the face of interference [56, 58]. Our work represents a foundational step in building a robust reasoning engine, but true physical intelligence will require continued progress on all these fronts.

### v4

@@ V-A2 Performance on the PhysiCa-Bench Benchmark
The FSPA directly measures the fidelity of the CPM as a physics simulator by calculating the Mean Squared Error between its predicted final state and the ground-truth state from the simulation. Our model achieves a low FSPA of 0.08, indicating that its internal world model generates accurate, quantitative predictions. Baseline LLMs cannot be evaluated on this metric as they lack the architectural components to produce a structured state prediction, a key differentiator of our approach. This capacity for state prediction is a foundational element of classic world models in robotics and reinforcement learning [Ha2018WorldModels, Hafner2023DreamerV3], and our work successfully integrates this principle into the LLM paradigm.
The most compelling evidence for our framework’s success comes from the Causal Consistency Score (CCS). This strict metric requires the model to correctly answer both a factual question and its corresponding counterfactual counterpart. A high CCS is difficult to achieve through superficial pattern matching; it requires the model to understand the specific causal link between the variable that was changed (the intervention) and the change in the outcome. Our CWMI model achieves a remarkable CCS of 87.6%. In stark contrast, Llama 3 and GPT-4 score only 12.3% and 21.9%, respectively. Their scores, while better than random chance, suggest they are answering the questions independently, with no consistent underlying causal model to link the factual and counterfactual pairs. They may correctly guess that a steel ball falls faster than a rubber one, but they fail to do so with the consistency that would imply genuine understanding. This finding strongly supports our core thesis: the interventional training paradigm described in Section IV (Eq. 6) is exceptionally effective at forcing a model to move beyond correlation and learn the causal structure of the physical world. This aligns with broader calls in the AI community for a more central role for causality in building robust and generalizable intelligence [Kiciman2024Causal, Dasgupta2021Grounded].
@@ V-B Ablation Studies: Deconstructing the CWMI Framework
To understand the source of CWMI’s performance and validate our specific design choices, we conducted a series of ablation studies on PhysiCa-Bench. We systematically removed or altered key components of our framework and measured the impact on performance.
@@ V-B3 Validating the Architectural Separation
Finally, to justify our two-component architecture, we conducted an experiment where we discarded the CPM entirely and instead fine-tuned the end layers of the Llama 3 LLM directly on the physical reasoning tasks. This ”LLM Fine-tuning (No CPM)” approach is representative of standard domain adaptation techniques. Its performance is poor across the board, barely improving upon the zero-shot GPT-4 baseline and showing almost no causal consistency. This result strongly validates our architectural decision to separate linguistic parsing from physical simulation. A standard Transformer architecture, while a universal approximator in theory, lacks the specific inductive biases for modeling physical dynamics efficiently. Our CPM, while also Transformer-based, is specialized for this role. It operates on a structured state representation and is trained with a physics-oriented objective, making it far more effective and parameter-efficient for this class of problems. This finding echoes the principles of modularity and relational inductive biases [Battaglia2018GNNs, Mott2019MIC], suggesting that building complex AI systems may require composing specialized modules rather than relying on a single, monolithic model for all tasks.
@@ V-C1 The Effect of Causal Physics Module Capacity
This trend is highly informative. The initial sharp increase in performance suggests that a certain model capacity is required to capture the fundamental physical laws present in our dataset. As the model grows, it becomes more adept at modeling complex interactions and second-order effects. The eventual plateau indicates that, for the complexity of physics represented in PhysiCa-Bench (primarily rigid-body dynamics), our 256M-parameter CPM is near-optimal. This scaling analysis provides confidence in our model design and suggests a clear path for future work: to tackle more complex physical domains like fluid or non-rigid dynamics, a correspondingly larger and perhaps architecturally different CPM will be required. This mirrors the scaling laws that have been so fundamental to the progress of LLMs themselves [Brown2020GPT3].
@@ V-C2 Qualitative Analysis: When It Works and When It Fails
These limitations are not surprising; they define the boundaries of our current contribution and illuminate the path forward. Addressing them will likely require architectural enhancements to the CPM, such as adopting graph neural networks for multi-object tracking [Battaglia2018GNNs], incorporating physics-informed neural network techniques for continuous fields [Raissi2019PINNs], or integrating more sophisticated 3D visual representations [Lerer2019Pytorch3D]. Furthermore, creating more robust systems that can handle noisy or incomplete descriptions is a shared challenge across many AI domains, from facial expression recognition from heterogeneous sources to building reliable WiFi-based sensing systems in the face of interference [9613773]. Our work represents a foundational step in building a robust reasoning engine, but true physical intelligence will require continued progress on all these fronts.